A researcher collected the following information for two variables x and y

No. of pairs = 20, r = 0.5, mean of X is 15, mean of Y is 20, standard deviation of X is 4, 

standard deviation of Y is 5.

Later on it was found that one pair of value as been wrongly taken as (X = 16, Y = 30) whereas 

the correct values were (X = 26, Y = 35). Find the correct value of r.


1
Expert's answer
2020-08-12T16:38:05-0400

As per the question,

no. of pairs say n=20

Initially Coefficient of correlation r =0.5

mean X=15

mean Y=20

standard deviation for X say "s_x" =4

standard deviation for Y say "s_y" =5

Initially the pair was (15,20) for this pair only lets calculate "z_x" and "z_y"

"z_x" ="\frac {x_i-mean(X)}{s_x}"

Similarly,

"z_y"="\frac{y_i-mean(y)}{s_y}"

so "z_x"="\frac{16-15}{4}"

=0.25

similarly

"z_y" ="\frac{30-20}{5}"

=2

let calculate the product of "z_x" and "z_y" =say m=0.5

let the sum of the product of "z_x" and "z_y" of the remaining products be s.

using formula,

r="\frac{\sum(z_x\times z_y)}{n-1}"

0.5="\frac{s+0.5}{20-1}"

on solving we get s=9 .....(equation 1

Now the correct pairs was(26,35)

Similarly calculating the value of"z_x and z_y" using the above mention formula

we get,

"z_x" =2.75

Similarly

"z_y" =3

Now the product of "z_x and z_y" =2.75"\times" 3

=8.25

Now again using the formula for coefficient of correlation

r="\frac{\sum(z_x\times z_y)}{n-1}"

="\frac{(8.25+s)}{20-1}"

putting the value of s from equation 1 we get

r="\frac{8.25+9}{19}"

="\frac{17.25}{19}"

r=0.9078

Hence the correct value of r is 0.9078.